i do not understand this at ALL, PLEASE help. the question is: write an equation for each line and graph it.
through (2,5) and parallel to 2x+4y=8
i don't know how to change 2x+4y=8 into a form so that i can graph it.
the other one is:
through (2,5) and perpendicular to 2x+4y=8
i'm lost, please , please help me!!

Answer by Alan3354(69443)

through (2,5) and parallel to 2x+4y=8
i don't know how to change 2x+4y=8 into a form so that i can graph it.
2x + 4y = 8 is a straight line, so only 2 points are needed.
Change it to slope-intercept form y = mx + b (that means solve for y).
x + 2y = 4
2y = -x + 4
y = (-1/2)x + 2
The slope, m, = -1/2 and the y-intercept is b.
All lines parallel have the same slope.
Now, find the equation of the line thru (2,5) with a slope of -1/2.
Use y = mx+b
5 = (-1/2)*2 + b
b = 6
--> y = (-1/2)x + 6
or x + 2y = 12
--------------------
the other one is:
through (2,5) and perpendicular to 2x+4y=8
The slope of lines perpendicular is the negative inverse of -1/2, so
m = +2
5 = 2*2 + b
b = 1
y = 2x + 1 is perpendicular thru (2,5)
--------------
To graph these, find 2 points.
y = 2x+1
Pick 2 values for x and find y
x = 0, y = 1
x = 3, y = 7
(0,1) and (3,7)